# MedHalu: Hallucinations in Responses to Healthcare Queries by Large Language Models

### v1

@@ 4.4 Hallucination Detection per Hallucination Type
Next, we study hallucination detection for each of the hallucination types to check if LLMs can detect some hallucination types better than the others. Table 7 shows the results for each hallucination type. As we observed in Section 4.3, GPT-3.5 and GPT-4 perform better than LLaMA-2 in hallucination detection overall. Upon diving deeper into each of the hallucination types, LLaMA-2 can detect context-conflicting and fact-conflicting hallucinations better than the input-conflicting hallucination for HealthQA subset. On the contrary, it detects input-conflicting hallucination better for LiveQA with an average macro-F1 of , whereas the best macro-F1 for MedicationQA is also but for fact-conflicting hallucination type. GPT-3.5 and GPT-4 give clear indications of detecting context-conflicting hallucination the best for all the subsets, followed by fact-conflicting and input-conflicting hallucination types. Intuitively, it makes sense as well since it is easier to detect self-conflicts in context-conflicting hallucinations just by looking at the LLM-generated healthcare responses. Conversely, fact-conflicting hallucination is challenging since it demands prior medical knowledge to be able to detect the presence of fact-conflicts. Similarly, input-conflicts are also slightly difficult to detect since it requires detecting conflicts with the input system prompt and the healthcare query.
@@ 5 Expert-in-the-loop to Improve LLM Hallucination Detection
We observe in Section 4.3 that LLMs are slacking behind the experts by a very large margin in medical hallucination detection. On an average, they perform no better than laypeople and in some cases, even worse than them. Therefore, in this section, we explore the possibility to improve the detection of medical hallucinations using LLMs by leveraging expert reasoning.
We therefore go back to the human experts, and ask them to articulate their reasoning about how they detected hallucinations and the reason behind their decision of whether a medical response is hallucinated or not. Their responses indicated that a combination of prior domain knowledge as well as looking up trusted health-related public websites such as UpToDate, BMJBestPractice, WebMD and NHS (UK) were used to cross-verify the answers. Therefore, we propose expert-in-the-loop approach to enable experts to help in improving hallucination detection using LLMs. We basically feed in experts’ reasoning together with healthcare queries and generated responses into the LLMs and evaluate their hallucination detection performance. The hallucination detection prompt with expert-in-the-loop approach is shown in Table 8.
I want you to act as an answer judge. Given a medical question, answer and an expert reason, your objective is to detect if the answer contains non-factual or hallucinated information. You should give your judgment based on the following 3 hallucination types, expert reason and the world knowledge.
@@ 6.1 Large Language Models
Large Language Models (LLMs) such as GPT-4 [24], LLaMA-3.1 [25], Claude-3 [26], Mistral [27], and Gemini [9] have demonstrate significant instruction-following capabilities [28, 29] and achieved substantial success across diverse general-purpose language modeling tasks including classification [30, 31], reasoning [32, 33, 34], and summarization [35, 36, 37]. Their proficiency extends to handling complex medical inquiries by integrating expert knowledge and advanced reasoning abilities [38, 39, 40, 41]. However, their high proficiency can mislead users into overestimating their reliability [42, 43, 44], leading to trust in outputs that may be factually inaccurate [45, 46].
@@ 6.2 Hallucinations in LLMs
As large language models (LLMs) like ChatGPT become widely used in public domains, concerns about their tendency to generate hallucinated content have intensified [12, 47, 48, 49, 50, 51]. Hallucination in LLMs is defined as content that, while often appearing plausible, is nonsensical or unfaithful to the source and factually incorrect, thereby complicating detection efforts [21, 52]. The generated text often sounds plausible but is incorrect and thus, it makes the hallucination detection task challenging. [20] categorizes hallucinations into input-conflicting, context-conflicting and fact-conflicting which reflect deviations from user input, internal inconsistencies, and inaccuracies against established facts, respectively.
@@
Efforts to systematically evaluate these phenomena have led to the development of benchmarks such as HaluEval [53], which assesses hallucinations using three tasks, including question answering, knowledge-grounded dialogue, and text summarization. In the healthcare domain, the Medical Domain Hallucination Test (Med-HALT) [14] leverages a multinational dataset to test LLMs on medical multiple-choice questions, focusing on reasoning and memory-related hallucinations. [54] introduced UPHILL, a dataset of health-related claims that tests LLMs’ abilities to handle increasing levels of presuppositions and factual inaccuracies. Our study is the first to address hallucinations in responses to real-world healthcare queries from patients, an under-explored research field.
7 Conclusion and Future Works
We propose MedHalu, a first-of-its-kind LLM-generated medical hallucination dataset with medical queries from diverse health topics and the hallucinated LLM responses, coupled with the labeled hallucination types and text spans. Experiments with three groups of evaluators – medical experts, LLMs, and laypeople – reveal that LLMs still fall short in understanding and reasoning about complex, domain-specific information, leading to poor hallucination detection in medical domains.
The fact that LLMs perform worse than medical experts and, in some cases, no better or even worse than laypeople in detecting hallucinations raises concerns about their readiness for real-world applications where accuracy is paramount. This suggests that while LLMs can be powerful tools, they may introduce risks when used without proper oversight, particularly in contexts requiring specialized knowledge. Looking forward, we propose the following future directions:
@@ 7 Conclusion and Future Works
Enhancing Expert Feedback Loops: Building on our proposed expert-in-the-loop approach, future work could focus on refining mechanisms that allow LLMs to continuously learn from expert feedback. This could involve interactive systems where LLMs not only generate responses but also seek validation or corrections from experts in real-time;
@@
Extension to Multilingual and Multimodal Scenarios. While our study primarily focuses on English medical queries in textual formats, future research should explore the generalizability of the findings to non-English languages and LLM hallucination under alternative modalities, such as medical videos, to address language-specific and domain adaptation challenges.
Hybrid Approaches: Future works can develop hybrid approaches that combine LLMs with rule-based systems or knowledge graphs that encode expert knowledge. Such approaches could mitigate the risk of hallucinations by cross-referencing generated content with verified medical information;
Longitudinal Studies: Conducting longitudinal studies to track LLMs’ susceptibility to hallucinations over time, particularly as they are exposed to new data and contexts, will be crucial in understanding how these models evolve and whether their performance in detecting hallucinations improves.
8 Ethical Usage of Dataset
@@
9 Acknowledgements
Appendix A Prompts for Hallucination Generation

### v2

@@ 3.1 Methodology
Hallucination detection by Experts and Laypeople. We employ groups of medical experts and laypeople through Prolific. Medical experts are selected only if they are native English speakers based in the UK and have graduated with at least an undergraduate degree in health/medicine. Similarly, only those laypeople, who are native English speakers based in the UK but do not have any degree or background in healthcare/medicine, are selected. In order to keep the costs of human evaluation in check, we randomly sampled the MedHalu dataset using stratified sampling of the base datasets—HealthQA, LiveQA, and MedicationQA. We sample question answer pairs in total. We randomly split question-answer pairs into batches, each containing pairs. We hire evaluators for each batch to evaluate question-answer pairs in two hours. Therefore, we employ medical experts and laypeople in the overall study. The detailed annotation process is discussed in Section 2.3 and Appendix B.
3.2 Experimental Setup
For generating hallucinated responses to the healthcare queries, we use GPT-3.5 (Ouyang et al. 2022) using OpenAI’s official API. We set temperature to and maximum generation length to tokens. For detecting LLM hallucinations, we input our detection prompt into each of the LLMs together with the healthcare query and corresponding response. For LLaMA-2-Instruct (Touvron et al. 2023), we use its open-source implementation after downloading the weights for model with billion parameters. For OpenAI’s GPT-3.5 and GPT-4, we use their official API. We set the same temperature of and maximum generation length of for all the LLMs. Our primary goal in including models with different parameters is not to claim superiority of any one model but to provide context for interpreting results and testing the challenging MedHalu benchmark.
3.3 Evaluation Metrics
We model hallucination detection as a binary classification task and thus leverage accuracy, macro precision (ma-P), macro-Recall (ma-R), and macro-F1 scores (ma-F1) as the evaluation metrics. We also ask the evaluators to highlight text spans which contain hallucinations. To measure the effectiveness of detecting hallucinated text spans, we measure the edit distance between the LLM-detected text spans and the expert annotated hallucinated spans. Edit distance measures the minimum number of changes (insertion, deletion, or substitution of characters) required to convert one string into the other. Smaller edit distances indicate greater similarity between LLM detected text spans and expert annotations.
4 Results
4.1 Overall Results
Performances of LLMs. Table 6 shows the results of different evaluator groups in hallucination detection on MedHalu. For HealthQA subset, LLaMA-2 achieves an F1-score of whereas GPT-3.5/4 achieve higher scores of and , respectively. On the more challenging LiveQA dataset, which contains real consumer health queries received by the U.S. National Library of Medicine, GPT-3.5/4 both achieve F1-score of only . For MedicationQA, the highest F1-score is . Overall, the proprietary models GPT-3.5/4 significantly outperform the open source LLaMA-2 for hallucination detection, with GPT-4 showing only marginal improvement over GPT-3.5.
Performances of Human Experts and Laypeople. Medical experts achieve macro-F1 scores of for HealthQA, for LiveQA, and for MedicationQA. The consistently low accuracy and macro-F1 scores highlight the difficulty of hallucination detection in LiveQA even for trained professionals. As expected, laypeople perform much worse than the experts, achieving macro-F1 scores of only , and for the three datasets and therefore, are more vulnerable to these hallucinated healthcare responses. Surprisingly, LLMs perform no better than laypeople except in the LiveQA subset, indicating that LLMs struggle with hallucination detection on specialized domains due to the lack of domain knowledge and are even unable to detecting self-generated hallucinated responses to the healthcare queries.
Results for detecting hallucinated text spans. We next evaluate the capabilities of different LLMs for detecting hallucinated text spans. Correctly identifying these hallucinated text spans is important because: a) this can allow the LLMs to be improved to generate fewer hallucinations; b) it can help laypeople and experts ignore the hallucinated part of the text. We consider the expert annotated hallucinated text spans for question-answer pairs as ground truth. We then calculate the edit distance between all the possible combinations of LLM-detected and expert annotated text spans and select the minimum score for each ground-truth hallucinated text span. Table 7 shows the mean and median edit distance values between the LLM-detected and the expert-detected hallucinated text spans for each of the subsets in MedHalu dataset. We exclude LLaMA-2 because it was incapable of detecting hallucinated text spans during our initial experiments even though we tried with various different prompts. GPT-3.5 achieves mean edit distance values of , , and for HealthQA, LiveQA and MedicationQA, respectively. On the other hand, GPT-4 achieves mean edit distance values of , , and for HealthQA, LiveQA and MedicationQA, respectively. Out of GPT-3.5 and GPT-4 models, GPT-4 consistently has a higher agreement with expert evaluators as evident from its lower edit distance values. LiveQA gets the highest edit distance values among the three subsets for both GPT-3.5 and GPT-4, again indicating that LiveQA is a challenging dataset. For the benefit of the research community, we will also make these LLM detected hallucinated text spans publicly available to allow fine-grained hallucination detection.
4.2 Are some types of hallucination easier to detect than others?
@@
Next, we study hallucination detection for each of the hallucination types (Section 2.1) to check if LLMs can detect some hallucination types better than the others. Table 8 shows the results for each hallucination type. As we observed in Section 4, GPT-3.5 and GPT-4 perform better than LLaMA-2 in hallucination detection overall. Upon diving deeper into each of the hallucination types, LLaMA-2 can detect context-conflicting and fact-conflicting hallucinations better than the input-conflicting hallucination for HealthQA subset. On the contrary, it detects input-conflicting hallucination better for LiveQA with an average macro-F1 of , whereas the best macro-F1 for MedicationQA is also but for fact-conflicting hallucination type. GPT-3.5 and GPT-4 give clear indications of detecting context-conflicting hallucination the best for all the subsets, followed by fact-conflicting and input-conflicting hallucination types. This also makes sense since it is easier to detect self-conflicts in context-conflicting hallucinations just by looking at the LLM-generated healthcare responses. Conversely, fact-conflicting hallucination is challenging since it demands prior medical knowledge to be able to detect the presence of fact-conflicts. Similarly, input-conflicts are also slightly difficult to detect since it requires detecting conflicts with the input system prompt and the healthcare query.
@@ 5 Expert-in-the-loop to Improve LLM Hallucination Detection
We observe in Section 4 that LLMs underperform compared to experts by a very large margin in medical hallucination detection. On average, they perform no better than laypeople and in some cases, even worse than them. Therefore, in this section, we explore the possibility to improve the detection of medical hallucinations using LLMs by leveraging insights on how experts reason about the presence of hallucinations.
We therefore go back to the human experts, and ask them to articulate their reasoning about how they detected hallucinations and the reason behind their decision of whether a medical response is hallucinated or not. Their responses indicate that a combination of prior domain knowledge as well as looking up trusted health-related public websites such as UpToDate, BMJBestPractice, WebMD and NHS (UK) were used to cross-verify the answers. Therefore, we propose expert-in-the-loop approach to enable experts to help in improving hallucination detection using LLMs. We feed in experts’ reasoning together with healthcare queries and generated responses into the LLMs and evaluate their hallucination detection performance. The hallucination detection prompt with expert-in-the-loop approach is shown in Table 9.
I want you to act as an answer judge. Given a medical question, answer, and an expert reason, you should detect if the answer contains non-factual or hallucinated information. You should give your judgment based on the following 3 hallucination types, expert reasons, and world knowledge.
@@
You should try your best to determine if the answer contains non-factual or hallucinated information according to the above hallucination types. The answer must be either Yes or No. If Yes, provide the non-factual or hallucinated text spans from the answer in a bullet format without any other information.
@@ 6.1 Large Language Models
Large Language Models (LLMs) such as GPT-4 (OpenAI 2023b), LLaMA-3 (AI 2024), Claude-3 (Anthropic 2024), Mistral (Jiang et al. 2023), and Gemini (Team et al. 2023) have achieved substantial success across diverse general-purpose language modeling tasks including classification, reasoning, and summarization (Srivastava et al. 2023; Zhu et al. 2023; Liu et al. 2023; Jin et al. 2024a, b). Their proficiency extends to handling complex medical inquiries by integrating expert knowledge and advanced reasoning abilities (Nori et al. 2023; Singhal et al. 2023a, b; Liévin et al. 2023). However, their high proficiency can mislead users into overestimating their reliability, leading to trust in outputs that may be factually inaccurate (Chen et al. 2024b).
@@ 6.2 Hallucinations in LLMs
As LLMs become widely used in public domains, concerns about their tendency to generate hallucinated content have intensified (Rawte, Sheth, and Das 2023; Deng et al. 2024; Chen et al. 2024c). Hallucination in LLMs is defined as content that, while often appearing plausible, is nonsensical or unfaithful to the source and factually incorrect, thereby complicating detection efforts (Ji et al. 2023; Chen et al. 2023b, a; Liu et al. 2022; Xu, Jain, and Kankanhalli 2024; Zhao et al. 2024). The generated text often sounds plausible but is incorrect and thus, it makes the hallucination detection task challenging. Zhang et al. (2023) categorizes hallucinations into input-conflicting, context-conflicting and fact-conflicting which reflect deviations from user input, internal inconsistencies, and inaccuracies against established facts, respectively.
6.3 Benchmarks for LLM Hallucinations
@@
Efforts to systematically evaluate LLM hallucinations have led to the development of generic benchmarks such as HaluEval (Li et al. 2023), which assesses hallucinations in general contexts using three tasks, including question answering, knowledge-grounded dialogue, and text summarization. In the healthcare domain, the Medical Domain Hallucination Test (Med-HALT) (Pal, Umapathi, and Sankarasubbu 2023) provides a multinational dataset of multiple choice questions derived from medical examinations across various countries, focusing on reasoning and memory-related hallucinations. Vishwanath et al. (2024) detects hallucinations in the generated summaries from Electronic Health Records. Kaur, Choudhury, and Pruthi (2023) introduce UPHILL, a dataset of health-related claims that tests LLMs’ abilities to handle increasing levels of presuppositions and factual inaccuracies. All of these benchmarks are very different from the context of a lay person asking an LLM their medical queries, which is a highly common use case, especially after the prevalence of ChatGPT and other chatbots. Therefore, we introduce MedHalu and present the first study to address LLM hallucinations in responses to real-world healthcare queries from layperson patients.
7 Conclusions and Future Works
We propose MedHalu, a pioneering hallucination detection benchmark featuring diverse healthcare queries and corresponding LLM responses, annotated with hallucination types and text spans. Evaluation on medical experts, LLMs, and laypeople highlights that the MedHalu benchmark is challenging and also reveals the current limitations of LLMs in detecting hallucinations, particularly in complex, domain-specific scenarios.
This is precisely the motivation of our study, to highlight this current limitation, and to create a strong benchmark that can drive forward future research to overcome such limitations (e.g., through knowledge graphs (as noted below), medical literature and expert reasoning)
Looking forward, we propose several key directions for future research:
Mitigating Hallucination through Adaptation. MedHalu offers a rich corpus for fine-grained LLM hallucination detection. Fine-tuning LLMs using parameter-efficient techniques, such as LoRA (Hu et al. 2022) and QLoRA (Dettmers et al. 2024), on MedHalu can improve their reliability in real-world healthcare queries. Moreover, combining LLMs with rule-based systems, knowledge graphs (Agrawal et al. 2024) such as uncertainty-aware knowledge graphs (Chen, Chen, and Yang 2024) and medical literature that encode expert knowledge can mitigate hallucination risks by cross-referencing responses with verified medical information.
@@ 7 Conclusions and Future Works
Enhancing Expert Feedback Loops. Building on our proposed expert-in-the-loop approach, future work could focus on refining mechanisms that allow LLMs to continuously learn from expert feedback. This could involve interactive systems where LLMs not only generate responses but also seek validation or corrections from experts in real-time.
@@
Extension to Multilingual and Multimodal Scenarios. While our study primarily focuses on English medical queries in textual formats, future research can explore how LLMs handle inaccurate information in non-English languages and LLM hallucination under alternative modalities (e.g. medical videos and broadcast).
Limitations
The proposed MedHalu dataset contains real-world healthcare queries in English only. Therefore, it is unknown how LLMs would hallucinate in case of healthcare queries in non-English languages. In the future, we would like to focus on non-English queries as well to study LLM hallucinations. One possible approach can be to directly translate English healthcare queries into non-English languages to curate a multilingual dataset. Moreover, we are also in discussion with doctors in India for future work to potentially explore creating a benchmark for Hindi, Bengali and Marathi, which are among the 3 most spoken Indian languages.
As existing LLMs continue to train on more and more datasets and new LLMs keep releasing, hallucination detection may become increasingly challenging. It is important to keep up with that pace and continuously evaluate their ability to generate hallucinated text in order to ensure their safety and reliability. As healthcare requires its own domain knowledge, future work can build an agentic AI that can incorporate knowledge from either knowledge graphs or from specific sources of high quality medical literature (e.g., WebMD888https://www.webmd.com/ or Mayo Clinic articles999https://www.mayoclinic.org/diseases-conditions from the US, NHS help articles in the UK101010https://www.nhs.uk/health-a-to-z/ and others cited by our human experts as the strategies they use themselves in Section 5) to create better hallucination detectors for the limited context of real-world medical queries. This could also become part of upcoming AI regulations – to make high quality answers and/or the addition of hallucination detection filters as a key step for LLM providers.
Ethical Considerations
Ethical Usage of Dataset
@@
Longitudinal Studies. The fact that LLMs perform worse than medical experts and, in some cases, no better or even worse than laypeople in detecting hallucinations raises concerns about their readiness for real-world applications where accuracy is paramount. This suggests that while LLMs can be powerful tools, they may introduce risks when used without proper oversight, particularly in contexts requiring specialized knowledge. Conducting longitudinal studies to track LLMs’ susceptibility to hallucinations over time, particularly as they are exposed to new data and contexts, will be crucial in understanding how these models evolve and whether their performance in detecting hallucinations improves.
Appendix A Hallucination Generation Prompts
@@ A.3 Context-conflicting Hallucination
Appendix B Annotation Platform For Human Evaluation
For LLM hallucination detection, we hire two sets of human evaluators–medical experts and laypeople through Prolific. We develop a customized annotation platform for annotating LLM hallucinated responses to the healthcare queries. The screenshot of the annotation guidelines page is shown in Figure 1. Figure 2 shows example annotation pages within the annotation platform and the set of questions asked to the evaluators in case they find the provided LLM generated answer to be hallucinated (Figure 2(a)) or correct (Figure 2(b)).